Part1: Meir, Benson, and Lau are partners and share income and loss in a 3:2:5 ratio (in percents: Meir, 30%; Benson, 20%; and Lau, 50%). The partnership's capital balances are as follows: Meir $168,000, Benson $138,000, and Lau $294,000. Benson decides to withdraw from the partnership. Prepare journal entries to record Benson's February 1 withdrawal under each separate assumption:

a. Benson sells her interest to North for $160,000 after North is approved as a partner.

b. Benson gives her interest to a son-in-law, Schmidt, and Schmidt is approved as a partner.

c. Benson is paid $138,000 in partnership cash for her equity.

d. Benson is paid $214,000 in partnership cash for her equity.

e. Benson is paid $30,000 in partnership cash plus equipment recorded on the partnership books at $70,000 less its accumulated depreciation of $23,200.

Part 2: Assume that Benson does not retire from the partnership described in Part 1. Instead, Rhode is admitted to the partnership on February 1 with a 25% equity. Prepare journal entries to record Rhode's entry into the partnership under each separate assumption: Rhode invests (a) $200,000 (b) $145,000 (c) $262,000.
Answer #1

Journal Entries

Part 1

  Debit Credit

a. Benson Capital $138,000

North Capital $138,000

b.Benson Capital $138,000

Schmidt Capital $138,000

c.Benson Capital $138,000

Cash $138,000

d.Benson Capital $158,000

Meir Capital $28,500

Lau Capital $47,500

Cash $214,000

e.Benson Capital $138,000

Accumulated Depreciation $23,200

- Equipment

Equipment $70,000

Cash $30,000

Meir Capital $22,950

Lau Capital $38,250

Calculations:

d.Meir's Capital = ($214,000 - $138,000)× 3/8 = $28,500

Lau's Capital = ($214,000 - $138,000) × 5/8 = $47,500

e.Meir's Capital = [$138,000 - ($70,000 - $23,200 + $30,000) × 3/8 = $22,950

Lau's Capital = [$138,000 - ($70,000 - $23,200 + $30,000) × 5/8 = $38,250

Part 2:

  Debit Credit

a.Cash $200,000

Rhode Capital $200,000

b.Cash $145,000

Meir Capital $12,375

Benson Capital $8,250

Lau Capital $20,625

Rhode Capital $186,250

c.Cash $262,000

Meir Capital $13,950

Benson Capital $9,300

Lau Capital $23,250

Rhode Capital $215,500

Explanation:

a.Rhode's Capital = ($168,000 + $138,000 + $294,000) = $600,000

($600,000 + $200,000) × 25% = $200,000

Since there is no difference, thus no bonus is paid.

b.Calculations:

($600,000 + $145,000) × 25% = $186,250

$145,000 - $186,250 = ($41,250)

Meir Capital = $41,250 × 3/10 = $12,375

Benson Capital = $41,250 × 2/10 = $8,250

Lau Capital = $41,250 × 5/10 = $20,625

Bonus is paid to new partner Rhode.

c.Calculations:

($600,000 + $262,000) × 25% = $215,500

$262,000 - $215,500 = $46,500

Meir Capital = $46,500 × 3/10 = $13,950

Benson Capital = $46,500 × 2/10 = $9,300.

Lau Capital = $46,500 × 5/10 = $23,250

Bonus is paid by new partner to old partners.